remove parenthesis and solve: show steps please
5(x+2)+3(2x+1)-(x-5)=10

Answer by jayanthihemkos@yahoo.com(87)

5(x+2)+3(2x+1)-(x-5)=10
5x+10+6x+3-x+5 = 10
(5x+6x-x)+(10+3+5)=10
10x+18=10
10x+18-18=10-18
10x = -8
10x/10x=-8/10
x = -0.8